OpenStudy (anonymous):

simplify 2log(3x)-3(log(x)-log(y-6))

OpenStudy (anonymous):

2log(3x)-3(log(x/(y-6))=log(3x)^2-log(x^3/(y-6)^3)= log((9x^2)/[x^3/(y-6)^3])=log([9x^2(y-6)^3]/x^3)= log([9(y-6)^3]/x)